OpenStudy (anonymous):

the orderd pairs (1,5), (2,25), (3,125), (4,625), (5,3125) represent a function. What is a rule that represnts this function A. y=x^5 B. y=5x C. y=5^x D. y=x+5

OpenStudy (anonymous):

in an ordered pair, the first number is the x value and the second one is y, try putting x and y values in each function and see which one is satisfied.

OpenStudy (anonymous):

K

OpenStudy (anonymous):

I think its a

OpenStudy (amoodarya):

try to find out relation between 5 , 25, 125, 625 ,3125 5 ,5^2 ,5^3 ,5^4 ,... so is 5^x

OpenStudy (anonymous):

its not A, try again. :) check at least 3 pairs....A is true for only first pair.

OpenStudy (anonymous):

The answern is c

OpenStudy (amoodarya):

i do not say (a) the answer is (c) because 5^1=5 5^2=25 5^3=125

OpenStudy (anonymous):

@amoodarya my reply was for @JhonDoe ...u answered it straightaway but i wanted to make him find out.

OpenStudy (anonymous):

Thanks

OpenStudy (amoodarya):